OpenStudy (anonymous):

e^x = 4.5 Choose... 0.2747 1.5041 0.2877 -1.2345

OpenStudy (anonymous):

take the natural log of both sides. ln(e^x) = ln(4.5) x = ln(4.5) <-- plug into a calculator.

OpenStudy (anonymous):

you must have a lot of these questions

OpenStudy (anonymous):

yea i have three left but they are to hard

OpenStudy (anonymous):

-3e^4x + 11 = 2 can u help me on this one too

OpenStudy (anonymous):

ok nuk i got 1.5040 is that right

OpenStudy (anonymous):

probably. the answer should be positive since ln(1) = 0. so anything above 1 should be positive. also for your second question, manipulate it such that e^4x is alone. tell me what you get, and i'll help you from there.

OpenStudy (anonymous):

so what u do to one side u do it to the other side right

OpenStudy (anonymous):

yep.

OpenStudy (anonymous):

i took out e in im left with 4x=4.5

OpenStudy (anonymous):

woah, wait. how did you take out the e? Take out everything but the e. -3e^4x + 11 = 2 -3e^4x = -9 e^4x = 3 ln(e^4x) = ln3 4x = ln3 x = (ln3)/4

OpenStudy (anonymous):

thanks i dont know how but i got the answer